OpenStudy (anonymous):

find dy/dx.. y= 3/8 (x) + 3/8 (sinxcosx) + 1/4 (cos^3 x sinx)

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }x + \frac{ 3 }{ 8 }sinxcosx + \frac{ 1 }{ 4 }\cos^3x sinx\]

OpenStudy (anonymous):

the final answer should be cos^4 x

hartnn (hartnn):

it is indeed cos^4 x still interested in finding how ?

OpenStudy (anonymous):

yep.. can u help me @hartnn

hartnn (hartnn):

sure, u know the product rule in differentiation ?

OpenStudy (anonymous):

yes..

hartnn (hartnn):

then tell me what u got the derivative of \(cos^3x sin x=?\)

OpenStudy (anonymous):

\[(\cos^3x) (\cos x) + (\sin x)(3)(\cos^2 x)(-\sin x)\]

hartnn (hartnn):

that is correct :) now u know the formula of sin 2x ?

OpenStudy (anonymous):

\[\sin 2x= 2sinxcosx\]

OpenStudy (anonymous):

am i right?

hartnn (hartnn):

yes! u see how we can use that to differentiate middle term easily... 3/8 sin x cos x = 3/16 sin2x after diff. 6/16 cos2x = 3/8 cos 2x ok with this ?

OpenStudy (anonymous):

where did u get 6/16 cos 2x??

hartnn (hartnn):

\(\huge \frac{d}{dx}sin2x=2cos2x\)

OpenStudy (anonymous):

ok.. and then.. what's next??

hartnn (hartnn):

the first term diff. is just 3/8. now the 'differentiation' part is over.Lets put it all together and see what we can simplify: \(\large (3/8)+3/8cos2x-(3/4)cos^2xsin^2x+(1/4)cos^4x\) now do u know the formula for cos2x ??

OpenStudy (anonymous):

there are 3 formulas for cos 2x

hartnn (hartnn):

right, lets get everything in terms of cos....so which formula has cos in it ?

OpenStudy (anonymous):

cos 2x= cos^2 x - sin^2 x cos 2x= 2cos^2 x - 1 these 2 formulas have cos

hartnn (hartnn):

lets use 2nd one, where we directly get \(1+cos2x=2cos^2x\) and let use the fact that \(sin^2x=1−cos^2x\) so now we have : \((3/4)cos^2x−(3/4)cos^2x(1−cos^2x)+(1/4)cos^4x\) can u simplify this and tell me what gets cancelled

OpenStudy (anonymous):

hmmm...can i cancell both (3/4)cos^2x ???

OpenStudy (anonymous):

hmm ok i'll get it ... cos^4 x will be left.

hartnn (hartnn):

but did u get how i got that step ? ask if there is any doubt in any step/term.

OpenStudy (anonymous):

yes.. I get it.. but about this.. sin^2 x = 1- cos^2 x where din u get it?/

hartnn (hartnn):

basic trignometric identity \(\huge \color{red}{sin^2x+cos^2x=1}\)

OpenStudy (anonymous):

ow... Thank You For your help :))

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

^^